Wersja testu A 25 września 2003r.

1. a. T b. T c. N d. T 2. a. N b. T c. T d. N 3. a. N b. T c. T d. T 4. a. N b. N c. N d. N 5. a. N b. T c. N d. T 6. a. T b. T c. N d. N 7. a. N b. T c. T d. N 8. a. N b. N c. T d. T 9. a. N b. T c. N d. N 10. a. N b. N c. T d. T 11. a. N b. N c. T d. N 12. a. T b. T c. N d. T 13. a. T b. T c. N d. N 14. a. T b. N c. N d. T 15. a. N b. N c. T d. N 16. a. N b. N c. T d. T 17. a. N b. N c. T d. T 18. a. N b. N c. T d. N 19. a. T b. N c. N d. N 20. a. T b. N c. N d. T